5. x2+4x+3
6. x2+6x-2
10. 3x2+4x-5
14. -4x2+8x+1

Answer by stanbon(75887)

5. x^2+4x+3
Find two numbers whose product is 3 and whose sum is 4
The numbers are 3 and 1
Put 3x+x in place of 4x to get:
x^2+3x+x+3
Factor the 1st two and the last two terms separately to get:
x(x+3)+1(x+3)
Factor out the common factor of (x+3) to get:
(x+3)(x+1)
6. x^2+6x-2
Use the quadratic formula to get:
x=[-6+-sqrt(36-4*-2)]/2
x=[-6+-sqrt44)]/2
x=[-6+-2sqrt11]/2
x=[-3+-sqrt11]
Factors are
(x+3+sqrt11)(x+3-sqrt11)
----------------
10. 3x^2+4x-5
x=[4+-sqrt(16-4*3*-5)]/6
x=[4+-sqrt(76)]/6
x=[4+-2sqrt19)/6
x=[2+-sqrt19]/3
x=(2/3)+(1/3)sqrt19) or x=(2/3)-(1/3)sqrt19
Factors: (x-[(2/3)+(1/3)sqrt19)][x-[(2/3)-(1/3)sqrt19]]
14. -4x^2+8x+1
x=[-8+-sqrt(64-4*-4*1)]/-8
x=[-8+-sqrt80]/-8
x=[1+-(1/2)sqrt5]
Factors: (x-[1+(1/2)sqrt5]][x-[1-(1/2)sqrt5]]
--------------
Cheers,
Stan H.